The perimeter of a rectangle is 38 cm. the length is 11 cm longer than the width. Find the dimensions of the rectangle.
Answer by sofiyac(983)

let width be x
then length is x+11
2x+2x+22=38
4x+22=38
4x=16
x=4
so width is 4, then length is 4+11=15